I am having trouble solving the proportion for: x-4 divided by 5 = 2x + 3 divided by 9

Answer by arunpaul(104)

((x-4)/5)= ((2x+3)/9)
or ((x-4)*9) = ((2x+3)*5)
or 9x-36 = 10x +15
or 9x-10x =15+36
or -x = 51
or x =-51 (Ans)